OpenStudy (anonymous):

It takes a car 1.75 h to go from mile marker 10 to mile marker 115. What is the average speed of the car? A. 1.6 mph B. 60 mph C. 65 mph D. 183 mph

OpenStudy (anonymous):

Please help i don't know how to do this!

OpenStudy (anonymous):

Its easy, time=1.75 hours distance=115-10 S=105miles S=vt v=S/t v=105/1.75 v=60mph Option B is Correct